Write an expression for the area of an isosceles triangle and simplify. The base is 1/2Y and the height is 2/5y + 4
Answer by venugopalramana(3286)

Write an expression for the area of an isosceles triangle and simplify. The base is 1/2Y and the height is 2/5y + 4
AREA=BASE*HT/2=(Y/2)(2Y/5 +4)/2=Y(2Y+5)/(2*2)=(2Y^2+5Y)/4